if a geometric sequence is 80, -40 , 20 . what are the next three terms ?
Answer by jim_thompson5910(35256)

You're dividing by -2 each time. So the next three terms are

20/(-2) = -10

-10/(-2) = 5

5/(-2) = -2.5
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------